Given P(E) = 0.39, P(F) = 0.15, and P(E and F) = 0.03, what is P(E or F)?

Answer by ikleyn(53751)

.

For P(E or F),  use a standard formula of probability of the UNION of events

    P(E or F) = P(E) + P(F) - P(E and F) = 0.39 + 0.15 - 0.03 = 0.51.    ANSWER

Solved.